# Hybrid deep learning and iterative methods for accelerated solutions of viscous incompressible flow

### v1

@@ 4 Conclusion
    Exhibits super-resolution capability, achieving accurate solutions on a $512\times 512$ grid despite being trained on $128\times 128$ data (16×16$\times$16 × resolution increase).
@@
The prototype implementation of HyDEA combines heterogeneous frameworks: $\mathtt{Pytorch}$ ($\mathtt{Python}$) for DLSM module and $\mathtt{PETSc}$ (C) for CG-type solvers, inducing GPU-CPU communication overhead. Additionally, the CNN-based DeepONet restricts applications to uniform Cartesian grids. Future efforts will focus on code optimization by minimizing data transfer latency between frameworks and extending HyDEA to non-uniform grids for broader applicability. HyDEA’s hybrid paradigm bridges the gap between data-driven efficiency and numerical robustness, offering a scalable pathway for high-fidelity flow simulations.
@@ Appendix A Analysis of convergence behavior of the CG method
Following the formulation of Eq. (49), the iteration residual vector at $k$th iteration can be expressed in the following discrete form:
@@ Appendix B Sensitivity analysis of neural network parameters $A$, $B$, $C$ and $T$ on HyDEA’s performance
Figure 33: The iterative residuals of solving the PPE for Cases III and V. (a) $10th$ time step. (b) $1000th$ time step. (c) $8000th$ time step.
The computational times for Cases I to IV are comparable, while Case V demonstrates notably inferior performance. This discrepancy likely stems from the increased number of network parameters, which hinders convergence to optimal network parameters within the limited number of training epochs. This limitation forces HyDEA to perform additional iterations to achieve the prescribed residual threshold, as demonstrated by the convergence curves in Fig. 33. Specifically, for case III, it takes $2$ rounds with $10$ iterations at $10\Delta t$, $1$ round with $4$ iterations at $100\Delta t$, and $1$ round with $2$ iterations at $8000\Delta t$; for case V, it takes $3$ rounds with $11$ iterations at $10\Delta t$, $2$ rounds with $6$ iterations at $100\Delta t$, and $2$ rounds with $6$ iterations at $8000\Delta t$.
@@
The results demonstrate that cases with $T=20$ and $30$ require higher computational time compared to $T=40$. The inferior performance of DeepONet at these lower channel numbers suggests that reduced feature space dimensionality compromises the model’s ability to accurately learn the functional mapping between $e_{k}$ and $r_{k}$ spaces. This limitation forces HyDEA to perform additional iterations to achieve the prescribed residual threshold, as evidenced by the iterative residuals of $T=20$ and $40$ in Fig. 35, ultimately increasing total computational time. Specifically, for $T=20$, it takes $3$ rounds with $11$ iterations at $10\Delta t$, $2$ rounds with $6$ iterations at $100\Delta t$, and $2$ rounds with $6$ iterations at $8000\Delta t$; for $T=40$, it takes $2$ rounds with $10$ iterations at $10\Delta t$, $1$ round with $4$ iterations at $100\Delta t$, and 1 round with 2 iterations at $8000\Delta t$.
Figure 35: The iterative residuals of solving the PPE for $T=20$ and $T=40$. (a) $10th$ time step. (b) $1000th$ time step. (c) $8000th$ time step.
@@ Appendix D Sensitivity analysis of $Num_{CG-type}$ and $Num_{DLSM}$ on HyDEA’s performance
    2D flow around an inline oscillating cylinder at $Re=100$ with grid resolution of $192\times 192$, which maintains inherently unsteady characteristics. The dataset is constructed using $m=7000$ and $b=0.6$

### v2

@@ 4 Conclusion
    Exhibits super-resolution capability, achieving accurate solutions on a $512\times 512$ grid despite being trained on $128\times 128$ data ($16\times$ resolution increase).
@@
The prototype implementation of HyDEA combines heterogeneous frameworks: $\mathtt{PyTorch}$ ($\mathtt{Python}$) for DLSM module and $\mathtt{PETSc}$ (C) for CG-type solvers, which incurs computational overhead. Moreover, the loss function that incorporates the coefficient matrix leads to prohibitively high computational cost for large-scale simulations. Additionally, the CNN-based DeepONet restricts its applications to uniform Cartesian grids. Future efforts will focus on several key directions: code optimization to enhance the computational efficiency, extending HyDEA to non-uniform grids and parameterized linear systems for broader applicability, refining the dataset construction method for better compatibility with various traditional iterative solvers, and incorporating domain decomposition concepts to enable large-scale simulations. HyDEA’s hybrid paradigm bridges the gap between data-driven efficiency and numerical robustness, offering a scalable pathway for high-fidelity flow simulations.
@@ Appendix A Analysis of convergence behavior of the CG method
Following the formulation of Eq. (52), the iterative residual vector at $k$th iteration can be expressed in the following discrete form:
@@ Appendix B Sensitivity analysis of neural network parameters $A$, $B$, $C$ and $T$ on HyDEA’s performance
Figure 36: Iterative residuals of solving the PPE for Cases III and V. (a) $10th$ time step. (b) $1000th$ time step. (c) $8000th$ time step.
The computational times for Cases I to IV are comparable, while Case V demonstrates notably inferior performance. This discrepancy likely stems from the increased number of network parameters, which hinders convergence to optimal network parameters within the limited number of training epochs. This limitation forces HyDEA to perform additional iterations to achieve the prescribed residual threshold, as demonstrated by the convergence curves in Fig. 36. Specifically, for case III, it takes $2$ rounds with $10$ iterations at $0\Delta t, 11$ round with $4$ iterations at $1000\Delta t$, and $1$ round with $2$ iterations at $8000\Delta t$; for case V, it takes $3$ rounds with $11$ iterations at $10\Delta t$, $2$ rounds with $6$ iterations at $1000\Delta t$, and $2$ rounds with $6$ iterations at $8000\Delta t$.
@@
The results demonstrate that cases with $T=20$ and $30$ require higher computational time compared to $T=40$. The inferior performance of DeepONet at these lower channel numbers suggests that reduced feature space dimensionality compromises the model’s ability to accurately learn the functional mapping from $r_{k}$ to $e_{k}$ spaces. This limitation forces HyDEA to perform additional iterations to achieve the prescribed residual threshold, as evidenced by the iterative residuals of $T=20$ and $40$ in Fig. 38, ultimately increasing total computational time. Specifically, for $T=20$, it takes $3$ rounds with $11$ iterations at $10\Delta t$, $2$ rounds with $6$ iterations at $1000\Delta t$, and $2$ rounds with $6$ iterations at $8000\Delta t$; for $T=40$, it takes $2$ rounds with $10$ iterations at $0\Delta t, 11$ round with $4$ iterations at $1000\Delta t$, and 1 round with 2 iterations at $8000\Delta t$.
Figure 38: Iterative residuals of solving the PPE for $T=20$ and $T=40$. (a) $10th$ time step. (b) $1000th$ time step. (c) $8000th$ time step.
@@ Appendix D Sensitivity analysis of $Num_{CG-type}$ and $Num_{DLSM}$ on HyDEA’s performance
    2D flow around an inline oscillating cylinder at $Re=100$ with grid resolution of $192\times 192$, which maintains inherently unsteady characteristics. The dataset is constructed using $m=7000$ and $b=0.6$.
@@
Appendix E Sensitivity analysis of iterative termination criterion on HyDEA’s performance
This section presents a sensitivity analysis to evaluate the impact of iterative termination criterion on the performance of HyDEA. The analysis is conducted based on the flow configuration described in Section 3.1.1, and more specifically:
  • 1.
    2D lid-driven cavity flow at $Re=1000$ using grid resolution of $128\times 128$, with $Num_{CG-type}=3$ and $Num_{DLSM}=2$.
To avoid over-extensive comparisons, the HyDEA (MGPCG+DLSM-1) are employed in the forecasting stage.
We analyze the impact of the iterative termination criterion expressed in $L2$-norm on the performance of HyDEA by evaluating three different thresholds: $\epsilon=10^{-5}$, $\epsilon=10^{-6}$ and $\epsilon=10^{-7}$. The iterative residuals of solving the PPE using MGPCG and HyDEA (MGPCG+DLSM-1) at three representative time steps are presented in Fig. 44.
Figure 44: Iterative residuals of solving the PPE under different iterative termination criteria at the $10th$, $100th$ and $1000th$ time steps. (a)-(c) $\epsilon=10^{-5}$. (d)-(f) $\epsilon=10^{-6}$. (g)-(i) $\epsilon=10^{-7}$.
The iterative residuals of solving the PPE exhibits noticeable discrepancies in later time steps under different iterative termination criteria. These differences stem from the fact that the initial guess for the PPE in each new time step is inherited from the previous step’s solution. Notably, HyDEA requires fewer iterations than MGPCG in all cases, demonstrating its superior performance.